# Preprocessing Methods for Memristive Reservoir Computing for Image Recognition

### v1

@@ II-C Reservoir Computing with Memristors
Implementing a multi-layer perceptron (MLP) with nonvolatile memristors as weights requires flattening the input image. In comparison, the DFN approach is more effective than a fully connected MLP because it avoids flattening the input image by using spatial-temporal processing, thereby reducing both the number of required memristors and the effective area. Moreover, since only the readout layer requires training, the energy and time for training are substantially lower. For example, if the input image is of size $28\times 28$ and an MLP with two hidden layers of $20$ neurons each and an output layer with $10$ neurons will require $28028^{2}\times 20+20\times 20+20\times 10=16,$ 280 trainable parameters and the same number of memristors to represent these weights. The same operation can be performed using RC by $3652$ memristors out of which $332$ do not need to be trained, as we will show further.
@@ III Preprocessing Methods
Before sending the input image into the reservoir, the image is converted into a series of spatio-temporal write pulses. These pulses are written sequentially into the volatile memristors that constitute the reservoir. In this paper, we evaluate and compare various preprocessing methods using the MNIST handwritten dataset, which consists of $60,000$ training images and $10,000$ testing images. Each image is $28\times 28$ pixels, with greyscale values ranging from 00 to $255$. We binarize these values by labeling any pixel above $25$ as ${}^{\prime}1^{\prime}$ (white) and any pixel $25$ or below as ${}^{\prime}0^{\prime}$ (black) to preserve the form of the digit. In the reservoir, a ${}^{\prime}1^{\prime}$ translates to a write voltage pulse to the memristor, while ${}^{\prime}0^{\prime}$ corresponds to no voltage pulse [10].
@@ III-A One Dimensional Data Input (1D)
Figure 3: Schematic of one-dimensional preprocessing. Each horizontal row of pixels is converted into a pulse train of write voltages and is sequentially injected into a volatile memristor.
@@
In this method, each horizontal row of pixels in the image is converted into a pulse train of write voltages. Each pulse train is sequentially injected into a volatile memristor (see Fig. 3). If the input image is of size $n\times m$, where $n$ is the number of rows and $m$ is the number of columns, the reservoir will have $n$ volatile memristors, and each pulse train has $m$ pulses. In the case of MNIST, since the images have 28 rows, the reservoir has 28 volatile memristors. Each volatile memristor accumulates the write pulses corresponding to ${}^{\prime}1^{\prime}$s in a given row; its volatile nature naturally decays between pulses, capturing not just the sum of ${}^{\prime}1^{\prime}$s but also their temporal distribution. The decay rate and the frequency of these input pulses must be calibrated to ensure an accurate row representation. However, using 1D input alone can limit classification performance because a single memristor may be overloaded by the entire row, reducing its ability to distinguish nuanced spatial features.
III-B Two-Dimensional Data Input (2D)
In the 2D approach, both the horizontal rows and vertical columns of the image are transformed into pulse trains of write voltage, as shown in Fig. 4. As in 1D data input, each pulse train is injected into a single volatile memristor. For a 2D input of size $n\times m$, the reservoir requires $n+m$ memristors. For MNIST $(28\times 28)$, this results in a reservoir size of $56$. Because the image is provided to the reservoir from two orthogonal directions, the system gains a richer spatial representation, boosting accuracy. However, the trade-off is a larger reservoir (double that of 1D). Furthermore, 2D preprocessing alone, like its 1D counterpart, can be insufficient for classification if each memristor stores too many pulses. Hence, combining it with input sectioning, as described next, ensures that each memristor receives a more manageable portion of the data.
Figure 4: Schematic of two-dimensional preprocessing. Each horizontal row and vertical column of the image is converted into pulse trains of write voltages and sequentially applied to the reservoir’s volatile memristors.
Figure 5: Schematic of input sectioning. Only a portion of each row is given to each memristor. In the image shown, $4$ sections are used and seven pixels are applied to one memristor, improving accuracy by distributing the image representation across multiple devices.
III-C Sectioning Input Rows
@@
In input sectioning, each row and/or column of the image is divided into multiple smaller segments, as shown in Fig. 5. Each segment (section) of pixels is converted into a shorter pulse train that is sent to a separate volatile memristor, making it easier for the memristor to represent the data accurately. For an input image of size $n\times m$ divided into $k$ sections, the reservoir size is $n\times k$ for 1D processing and $(n+m)\times k$ for 2D (Table I). For example, a $28\times 28$ MNIST image with four sections per row in 1D mode requires $28\times 4=112$ memristors, while 2D mode with six sections per row and column requires $(28+28)\times 6=336$ memristors. By reducing the temporal load on each memristor, sectioning significantly enhances the reservoir’s ability to capture intricate spatial features and thus improves classification performance.
@@ IV Proposed Preprocessing Method - Parity
Figure 6: Schematic of parity preprocessing with 1D preprocessing. As in 1D, each horizontal row is converted into a pulse train of write voltages. In addition, the parity operation — an XOR between adjacent rows — produces an image with the outline of the digit. This is also converted into pulse trains of write voltages and applied to separate memristors.
The two-dimensional properties of the images are important but are not fully exploited without further preprocessing. The proposed parity preprocessing method can enhance the network’s accuracy by highlighting various features appearing in consecutive rows. Specifically, it performs an XOR operation on the $i$-th and $(i+1)$-th rows, generating additional ”parity rows”. Because parity complements existing methods, it is used together with the 1D or 2D input – as well as the input section – to ensure a more faithful representation of the input data (see Fig. 6). For instance, in 1D $+$ parity, the original rows are converted to write-voltage pulse trains, and the XORed rows are injected into separate memristors. With parity enabled, the 1D reservoir requires $12n-$ 1 volatile memristors, and the 2D case requires $(n+m+n-1)$. For MNIST $(28\times 28)$, using four sections in 1D $+$ parity results in $(28+27)\times 4=220$ memristors, while 2D $+$ parity with six sections requires $(28+28+27)\times 6=498$ memristors. This method proves effective because the parity rows provide a sparse outline of the digits, helping to discern edges more clearly and thus improving classification performance. More sophisticated cross-row and column operations could be explored in future work, but here we employ the parity method because of its simplicity.
@@ V Evaluation
Before being converted into write pulses, each MNIST image is binarized and then transformed according to the chosen preprocessing method. For 1D preprocessing, the image remains in its original $28\times 28$ form. In 2D preprocessing, the image is rotated by $90\degree$ and appended below the original, producing a $56\times 28$ matrix. Similarly, in parity preprocessing, the XOR operation is performed on adjacent rows, and the resulting matrix is appended below the original image, extending the matrix by an additional 27 rows. Finally, the rows are divided into sections, creating the pulse trains to be applied to the volatile memristors.
@@
These read currents are rescaled and multiplied by the weight matrix of the readout layer. The readout layer was implemented as a logistic regression classifier, applying a sigmoid activation element-wise to the result of the matrix-vector product between the reservoir states and the weight matrix of dimensions $[N\times 10]$, where N is the number of reservoir states, and 10 is the MNIST digit classes. This matrix is randomly initialized. During training, the predicted output for each input sample was computed and compared to a one-hot encoded label. The prediction error, equivalent to the gradient of the binary cross-entropy loss, was used to update the weights via stochastic gradient descent (SGD) using the outer product of the input vector and the error. This process was repeated for $500$ epochs with a learning rate of $0.02$.
Only the readout-layer weight matrix is updated during training — following the standard forward pass, error calculation, and gradient-based backpropagation. After training, we measure the accuracy of the 10,000-image test set. The network was implemented and trained in MATLAB. In the following, we evaluate RC using the accuracy, throughput, energy efficiency, and area metrics and show the trade-offs among them.
@@ V-A Accuracy
Figure 7: Accuracy for different preprocessing methods.
Fig. 7 summarizes the test accuracy for the various preprocessing methods. 2D preprocessing increases the accuracy by approximately 7% (2%) compared to 1D without (with) parity. Input sectioning proves essential in boosting accuracy, as each memristor receives a smaller, more manageable portion of the data. Without sectioning, the accuracy remains below 75%. In 1D mode, adding more sections improves accuracy, but not sufficiently to match 2D or parity-enhanced configurations, whereas in 2D mode, excessive sectioning can over-fragment the representation and slightly reduce accuracy. Parity preprocessing yields a substantial accuracy gain — approximately 6% for 1D and 2% for 2D — by highlighting edges and offering richer input representation.
@@ V-B Throughput, Energy Efficiency, and Area
Here, throughput is determined as the number of images processed per second. The throughput grows almost linearly with the number of sections; it is independent of the number of dimensions and the usage of parity.
The energy consumption of reservoir memristors is dominated by write operations (over $94\%$), as each memristor is read only once per image, and the read voltage ($0.6V$) is significantly lower than the write pulse voltage ($1.5V$). Energy consumed per image increases approximately linearly with the number of dimensions and is roughly independent of the number of sections. The parity method increases energy consumption sub-linearly, as parity-related rows contain fewer ’1’s than the original data, leading to fewer write operations. We define energy efficiency as the number of images processed per joule.
The relative area is approximated by the number of RC memristors. Each memristor also includes the area of its input line (pulse train) and the output lines to the perceptron readout. The area grows roughly linearly with the number of dimensions, the use of parity, and the number of sections.
@@ Acknowledgment
This research has been partially funded by the European Union’s Horizon 2020 Research And Innovation Programme FET-Open NEU-Chip under grant agreement No. 964877.

### v3

@@ II-C Reservoir Computing with Memristors
Implementing a multi-layer perceptron (MLP) with nonvolatile memristors as weights requires flattening the input image. In comparison, the DFN approach is more effective than a fully connected MLP because it avoids flattening the input image by using spatial-temporal processing, thereby reducing both the number of required memristors and the effective area. Moreover, since only the readout layer requires training, the energy and time for training are substantially lower. For example, if the input image is of size 28$\times$28 and an MLP with two hidden layers of 20 neurons each and an output layer with 10 neurons will require $28^{2}\times 20+20\times 20+20\times 10=16,280$ trainable parameters and the same number of memristors to represent these weights. The same operation can be performed using RC by 3652 memristors out of which 332 do not need to be trained, as we will show further.
Prior relevant work includes [10], where varying input rate leads to different final memristor states, and [23], which employs a reservoir built using kernels comprising two non-volatile memristors and a capacitor producing two responses to the same input. The capacitor sets the decay rate but adds area and energy overhead.
@@ III Preprocessing Methods
Figure 3: Schematics of preprocessing methods. (a) One-dimensional preprocessing: each horizontal row of pixels is converted into a pulse train of write voltages and is sequentially injected into a volatile memristor. (b) Input sectioning: only a portion of each row is given to each memristor. In the image shown, 444 sections are used and seven pixels are applied to one memristor, improving accuracy by distributing the image representation across multiple devices. (c) Two-dimensional preprocessing: each horizontal row and vertical column of the image is converted into pulse trains of write voltages and sequentially applied to the reservoir’s volatile memristors. (d) Parity preprocessing with 1D preprocessing: as in 1D, each horizontal row is converted into a pulse train of write voltages. In addition, the parity operation — an XOR between adjacent rows — produces an image with the outline of the digit. This image is also converted into pulse trains of write voltages and applied to separate memristors.
Before entering the reservoir, each input image is converted into a sequence of spatio-temporal write pulses. These pulses are applied sequentially to the volatile memristors comprising the reservoir. We evaluate various preprocessing methods using the MNIST dataset, which includes 60,000 training and 10,000 test images. Each image comprises 28$\times$28 pixels with greyscale values from 0 to 255. These are binarized by labeling pixels above 25 as ’1’ (white) and others as ’0’ (black) to retain digit shape. In the reservoir, ’1’ translates to a write voltage pulse; ’0’ means no pulse [10].
@@ III-A One Dimensional Data Input (1D)
In this method, each horizontal row of pixels in the image is converted into a pulse train of write voltages. Each pulse train is sequentially injected into a volatile memristor (see Fig. 3(a)). If the input image is of size $n\times m$, where $n$ is the number of rows and $m$ is the number of columns, the reservoir will have $n$ volatile memristors, and each pulse train has $m$ pulses. In the case of MNIST, since the images have 28 rows, the reservoir has 28 volatile memristors. Each volatile memristor accumulates the write pulses corresponding to ${}^{\prime}1^{\prime}$s in a given row; its volatile nature naturally decays between pulses, capturing not just the sum of ${}^{\prime}1^{\prime}$s but also their temporal distribution. The decay rate and the frequency of these input pulses must be calibrated to ensure an accurate row representation. However, using 1D input alone can limit classification performance because a single memristor may be overloaded by the entire row, reducing its ability to distinguish nuanced spatial features.
III-B Sectioning Input Rows
@@
In input sectioning, each row and/or column of the image is divided into multiple smaller segments, as shown in Fig. 3(b). Each segment (section) of pixels is converted into a shorter pulse train and sent to a separate volatile memristor, improving accuracy by reducing the amount of data each memristor must represent. For an input image of size $n\times m$ divided into $k$ sections, the reservoir size is $n\times k$ for 1D processing and $(n+m)\times k$ for 2D (Table I). For example, a $\times 28$ 28 MNIST image with four sections per row in 1D mode requires 28×4=11228$\times$ 4=11228 × 4 = 112 memristors, while 2D mode with six sections per row and column requires (28+28)×6=336(28+28$\times$ 6=336( 28 + 28 ) × 6 = 336 memristors. By reducing each memristor’s temporal load, sectioning significantly enhances the reservoir’s ability to capture intricate spatial features and thus improves classification accuracy.
III-C Two-Dimensional Data Input (2D)
In the 2D approach, both horizontal rows and vertical columns of the image are converted into pulse trains of write voltage, as shown in Fig. 3(c). As in 1D data input, each pulse train is applied to a single volatile memristor. For a 2D input of size $n\times m$, the reservoir requires $n+m$ memristors. For MNIST (28×28)(28$\times$ 28)( 28 × 28 ), this results in a reservoir size of 565656. Because the image is applied from two orthogonal directions, the reservoir gains a richer spatial representation, boosting accuracy. The trade-off is a larger reservoir (double that of 1D). Furthermore, 2D preprocessing alone, like 1D, can be insufficient for classification if each memristor stores too many pulses. Hence, combining it with input sectioning ensures that each memristor receives a more manageable portion of the data.
@@ IV Proposed Preprocessing Method - Parity
[…]
@@ V Evaluation
Before being converted into write pulses, each MNIST image is binarized and then transformed according to the chosen preprocessing method. For 1D, the image remains in its original $\times 28$ 28 form. In 2D, the image is rotated by 90°90$\degree$90 ° and appended below the original, producing a 56×2856$\times$ 2856 × 28 matrix. Similarly, in parity preprocessing, the XOR operation is performed on adjacent rows, and the resulting matrix is appended below the original image, extending the matrix by 27 rows. Finally, the rows are divided into sections, creating the pulse trains to be applied to the volatile memristors.
@@
These read currents are first quantized to 6-bits, then rescaled using min-max scaling, and multiplied by the readout layers’ weight matrix. Quantization mimics an analog-to-digital converter (ADC) that may be used in hardware. It also helps assess the realistic accuracy that can be achieved in the presence of memristor variability. The readout layer was implemented as a logistic regression classifier, applying a sigmoid activation element-wise to the result of the matrix-vector product between the reservoir states and the weight matrix of dimensions $[N\times 10]$, where N is the number of reservoir states, and 10 is the MNIST digit classes. This matrix is randomly initialized. During training, the predicted output for each input sample was computed and compared to a one-hot encoded label. The prediction error, equivalent to the gradient of the binary cross-entropy loss, was used to update the weights via stochastic gradient descent (SGD) using the outer product of the input vector and the error. This process was repeated for 500500500 epochs with a learning rate of 0.020.020.02.
Only the readout-layer weight matrix is updated during training, using standard forward pass, error calculation, and gradient-based backpropagation. Accuracy is then measured on the 10,000-image test set. The network was implemented and trained in MATLAB. Memristor resolution was analyzed by quantizing the currents from 1-bit to 7-bits. Accuracy remains acceptable down to 4-bit quantization of memristor currents, but drops sharply below that. In the following, we evaluate RC using accuracy, throughput, energy efficiency, and area, highlighting trade-offs among these metrics.
@@ V-A Accuracy
Figure 4: Test accuracy across various preprocessing methods.
Fig. 4 summarizes the test accuracy across preprocessing methods. For the highest accuracy configuration (over 4 Sections), 2D improves the accuracy, on average, by 7% (3%) compared to 1D without (with) parity. Input sectioning proves essential in boosting accuracy, as each memristor receives a smaller, more manageable portion of the data. Without sectioning, the accuracy remains below 75%. In 1D mode, adding more sections improves accuracy, but not sufficiently to match 2D or parity-enhanced configurations, whereas in 2D mode, excessive sectioning can over-fragment the representation and slightly reduce accuracy. Parity preprocessing yields a substantial accuracy gain — approximately 6% for 1D and 2% for 2D — by highlighting edges and offering richer input representation. As compared to [10], we achieve 1-2% improvement in accuracy with fewer devices due to the parity method. While the accuracy is slightly lower than [23], our method offers the advantage of not requiring large capacitors that limit the scalability of memristive systems.
@@ V-B Throughput, Energy Efficiency, and Area
Here, throughput is determined as the number of images processed per second. The throughput grows almost linearly with the number of sections; it is independent of the number of dimensions and the usage of parity (see Fig. 5 (a)).
[…]
The relative area is approximated by the number of RC memristors. Each memristor also includes the area of its input line (pulse train) and the output lines to the perceptron readout. The area grows roughly linearly with number of dimensions, use of parity, and number of sections (see Fig. 5 (c)).
@@ Acknowledgment
This research has been partially funded by the European Union’s Horizon 2020 Research And Innovation Programme FET-Open NEU-Chip under grant agreement No. 964877 and funded by the European Union (ERC, Real-Database-PIM, 101157452). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.